OpenStudy (anonymous):

Which of the following is true? twenty-one is a composite number because one and twenty-one are its only factors. thirteen is a prime number because it is only divisible by one and itself. sixteen is a prime number because it is divisible by 4. twenty-eight is a composite number because eighty-two is a prime number.

OpenStudy (anonymous):

Which of the following is true? twenty-one is a composite number because one and twenty-one are its only factors. thirteen is a prime number because it is only divisible by one and itself. sixteen is a prime number because it is divisible by 4. twenty-eight is a composite number because eighty-two is a prime number.

OpenStudy (mathstudent55):

Whole numbers are either prime or composite. A prime number is a number that has exactly two factors, 1 and itself. A composite number has more than 2 factors.

OpenStudy (anonymous):

Thirteen is a prime number, because it's only factors are 1 and itself

OpenStudy (mathstudent55):

Look at the number 21. Can you list all its factors? What are all the numbers 21 is divisible by in addition to 1 and 21?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@mathstudent55 has a much better response than me, listen to him :)

OpenStudy (mathstudent55):

@1lorax2 Thanks of the compliment.

OpenStudy (anonymous):

Of course, I'm always ready to defer to somebody who's better equipped than myself to deal with a problem.

OpenStudy (mathstudent55):

Factors of 21: 1, 3, 7, 21. More than 2 factors. It is a composite number. Factors of 13: 1, 13. Exactly 2 factors. Prime number. Factors of 16: 1, 2, 4, 8, 16. More than 2 factors. Composite number. Factors of 28: 1, 2, 4, 7, 14, 28. More than 2 factors. Composite number.

OpenStudy (mathstudent55):

I have a question. Is 1 prime or not?

OpenStudy (anonymous):

What is the Least Common Denominator (LCD) of nine-sixteenths and nineteen twenty-fourths? forty-eight thirty-two sixteen 4

OpenStudy (anonymous):

If I remember right, I believe that 1 is a category all it's own.

OpenStudy (anonymous):

0 is composite, 1 is special, 2 is the only even prime, and so on and so forth

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

Correct. 1 is not prime because it does not have exactly 2 different factors. There is only one factor of 1, and it is 1.

OpenStudy (mathstudent55):

Now for the second question. The denominators are 16 and 24. You need to find the smallest number that is a multiple of both.

OpenStudy (anonymous):

To find the least common denominator, you have to factor both denominators. Then you choose the smallest factor that the two number have in common

OpenStudy (mathstudent55):

There are several ways of doing this. One way is list some multiples of both numbers until you find the firtt multiple that is common to both.

OpenStudy (mathstudent55):

To find multiples of a number, just multiply the number by 1, 2, 3, 4, ...

OpenStudy (anonymous):

Oh my gosh. I was completely wrong again. I don't even know what I was thinking of there- there's no operation that is similar. I completely defer to @mathstudent55 , so listen to him @swaggy_jackson

OpenStudy (mathstudent55):

Here are the first several multiples of 16 and of 24: 16, 32, 48, 64, 80, 96 24, 48, 72, 96, 120 Notice the smallest multiple they both have in common is 48. That means 48 is the LCD.

OpenStudy (mathstudent55):

This is one method. There are other methods. Another method of finding the LCD of two numbers is to first factor the two numbers into their prime factorizations.

OpenStudy (mathstudent55):

I'll use your problem as an example. We are looking for the LCD of 16 and 24, so we factor 16 and 24 into their prime factorizations.

OpenStudy (mathstudent55):

The prime factorizations are: \(16 = 2^4\) \(24 = 2^3 \times 3\)

OpenStudy (mathstudent55):

For the LCD, we take common factors with the highest exponent, and non-common factors. Common factors: 2 and 2. We need the higher exponent, so we pick \(2^4\). Non-common factors: there is only one, and it's 3, so we use it too. LCD = \(2^4 \times 3 = 16 \times 3 = 48\)